NEPTUNE Canada

News Details

July 2011 Mid-cruise Update

Jul 18, 2011

The first half of NEPTUNE Canada’s maintenance and installation cruise of 2011 has gone really well. We are currently working at Barkley Canyon, the fourth of the five node areas that we plan to visit and are well on our way to accomplishing our cruise objectives.

Corky the OctopusCorky the Octopus (Click to enlarge.)

Here is a summary from each of the node areas:

ODP 1027 @ 2660 metres (July 5 - 6)

NEPTUNE Canada (NC) first travelled to the ODP 1027 area to complete the following objectives:

  • Installation of the Ifremer-developed Piezometer used to study temperature and pressure changes in the ocean crust (see blog article for more details).
ROPOS connecting the Ifremer Piezometer.ROPOS connecting the Ifremer Piezometer. (Click to enlarge.)
  • Recovery of Bottom Pressure Recorders (BPRs) and of a Conductivity, Temperature and Depth sensor (CTD) that needed maintenance (calibration check), swapping it for a new CTD.
Bottom Pressure Recorder (BPR)Bottom Pressure Recorder (BPR) (Click to enlarge.)
Endeavour @ 2300 metres (July 7 - 11)

In September 2010, NC installed 2 of 3 trans-ridge cables that connect the node to study areas within the Endeavour Ridge Valley, specifically to the Main Endeavour Vent Field (MEF) and to the northern area of the ridge valley (RCMN location). The third node-valley cable will run south to Mothra Vent Field, and will be installed in September 2011.

Our first stop was at MEF where four instruments were installed in September 2010. They sent real-time data back to shore for a month, at which point we lost communications with the area. Such a problem could result, for example, from the failure of the junction box, cable, or a connector. Testing during this cruise has confirmed that the junction box works fine, which means the problem lies with the cable. Plans are to replace this cable in September 2011, which should then bring this suite of instruments online again (read blog).

Oily FlatOily Flat

While on-site at MEF, we also:

  • Recovered a CTD and the Benthic and Resistivity Sensor (BARS). BARS is used to monitor pure vent fluid properties, but had a hot vent growing over its cable (read blog).
Benthic and Resistivity Sensor (BARS)Benthic and Resistivity Sensor (BARS) (Click to enlarge.)
  • COVIS (acoustic imaging of hydrothermal vent flow/watch video) and RAS (water sampler) were fully tested (activated) following protocols given by PIs to ensure they were 100% operational. Both showed positive results and will be reactivated in September 2011 when communications are restored.
COVISCOVIS (Click to enlarge.)
  • Surveyed a site for the TEMPO-mini installation in September 2011. TEMPO-mini is a suite of instruments designed by Ifremer to study the vent animals and how they respond to changes in this extreme environment.

Next we headed to the Regional Circulation Mooring North (RCMN) location where 12 instruments by one mooring and a short period seismometer were installed in September 2010. During that cruise one cable failed and the cable originally intended to connect to Mothra was reallocated to RCMN. That was worthwhile as these instruments continue to provide quality real-time data to the scientific community. During this cruise we performed our first successful cable recovery. ROPOS cut the cable in half, then on separate dives, hooked into each portion of the cable and ascended to the surface. The cable pieces were then rolled onto the deck via the R.V. Thompson’s A-Frame. This difficult and delicate operation had to be done during calm daylight because while the ship was pulling up the cable it was tethered to the seafloor at 2300 metres deep!

ODP 889 @ 1250 metres (July 12)

Stopping at CORK (which NC will soon connect to, providing real-time access to the data), inspected the instrument platform, and deployed a CTD. We also recovered the Imagenex sonar, which was installed in May 2010 at Bubbly Gulch near the Bullseye Vent.

Barkley Canyon @ 400 - 1000 metres (July 13 - 14)

Our first stop was at the 400-metre upper slope area where we lost communication in February 2011 due to a possible strike by a trawler net. All instruments are present but most have been flipped over and will require advanced testing. There was extensive damage to the extension cable, which led to the loss of communication and will require a considerable repair effort (read blog). The team plans to retrieve all the instruments within a week so they can be tested and refurbished before an eventual redeployment next year.

Broken Cable at BarkleyBroken Cable at Barkley

We then visited the home of Wally the crawler, among the gas hydrate mounds on the wall of Barkley Canyon (870m). Wally was retrieved for cleaning and refurbishing, temperature probes were deployed, samples and experiments were collected, and a survey of the hydrate mound was carried out to document changes over recent months.

Finally, we recovered, maintained and redeployed Pods 1, 3 and 4 at the mid-canyon wall (900m), and canyon axis (1000m). Scientists are studying benthic (seafloor) ecology using suites of instruments on these and the Upper Slope Pod 2. Long-term monitoring will allow scientists to follow the seafloor response to pulses of food from the phytoplankton blooms of the surface waters, and observe changes in seafloor communities related to seasonal patterns of sediment transport and nutrient upwelling.

Pompom AnemonePompom Anemone
Sampling so far

Sampling is needed to calibrate and to extend the footprint of our network. We have collected 51 out of the planned 75 push cores samples, along with water samples (suction and Niskin bottles) for researchers during the cruise. At MEF, samples of tubeworms were taken for comparison of the symbiotic bacteria with other symbionts of tubeworms found elsewhere (read blog).

Tubeworm Sampling at a Black Smoker.Tubeworm Sampling at a Black Smoker. (Click to enlarge.)
Expectations for Cruise Part II:

We will remain at Barkley Canyon for the next while to focus on the recovery, inspection, maintenance and redeployment of instruments at all study areas there, including the upper slope at 400 meters, mid-canyon wall at 870-900 meters, and the canyon axis at about 1000 metres. The final stop will be maintenance at Folger Passage where we expect biofouling (things growing on our instruments) to be significant since it has been more than a year since our last visit to this one location in our network that receives sunlight. If time allows we will go back to Endeavour to do some surveying in preparation of the September cruise, and plan to dock in Esquimalt on July 26th.

The R/V Thompson departing Esquimalt. The crew is living on this ship for three weeks.The R/V Thompson departing Esquimalt. The crew is living on this ship for three weeks. (Click to enlarge.)

All ROPOS dives are recorded and can be found on SeaTube.


Return to News Home